# VAPO: Efficient and Reliable Reinforcement Learning for Advanced Reasoning Tasks

### v2

@@ Abstract
We present VAPO, Value-based Augmented Proximal Policy Optimization framework for reasoning models., a novel framework tailored for reasoning models within the value-based paradigm. Benchmarked the AIME 2024 dataset, VAPO, built on the Qwen 32B pre-trained model, attains a state-of-the-art score of $\mathbf{60.4}$. In direct comparison under identical experimental settings, VAPO outperforms the previously reported results of DeepSeek-R1-Zero-Qwen-32B and DAPO by more than 10 points. The training process of VAPO stands out for its stability and efficiency. It reaches state-of-the-art performance within a mere 5,000 steps. Moreover, across multiple independent runs, no training crashes occur, underscoring its reliability. This research delves into long chain-of-thought (long-CoT) reasoning using a value-based reinforcement learning framework. We pinpoint three key challenges that plague value-based methods: value model bias, the presence of heterogeneous sequence lengths, and the sparsity of reward signals. Through systematic design, VAPO offers an integrated solution that effectively alleviates these challenges, enabling enhanced performance in long-CoT reasoning tasks.
@@ 1 Introduction
In the Large Language Models (LLM) [3, 15, 11, 28, 25, 2, 4] RL training, value-free methods like GRPO [22] and DAPO [29] have demonstrated remarkable effectiveness. These approaches eliminate the computational overhead of learning a value model, instead computing advantage solely based on the final reward of the entire trajectory. The trajectory-level advantage is then directly assigned as the token-level advantage for each position in the sequence. When training a reliable value model is particularly challenging, value-free methods deliver an accurate and stable baseline for advantage calculation by averaging the rewards across multiple trajectories within a group. This group-based reward aggregation mitigates the need for explicit value estimation, which often suffers from instability in complex tasks. Consequently, value-free methods have gained significant traction in addressing difficult problems such as long-CoT reasoning, with substantial research efforts focused on optimizing their frameworks.
@@
Despite the notable success achieved by the value-free methods, we argue that value-based approaches possess a higher performance ceiling if the challenges in training value models can be addressed. First, value models enable more precise credit assignment by accurately tracing the impact of each action on subsequent returns, facilitating finer-grained optimization [21]. This is particularly critical for complex reasoning tasks, where subtle errors in individual steps often lead to catastrophic failures, and it remains challenging for model optimizing under value-free frameworks [30]. Secondly, in contrast to the advantage estimates derived from Monte Carlo methods in value-free approaches, value models can provide lower-variance value estimates for each token, thereby enhancing training stability. Furthermore, a well-trained value model exhibits inherent generalization capabilities, enabling more efficient utilization of samples encountered during online exploration. This significantly elevates the optimization ceiling of reinforcement learning algorithms. Consequently, despite the formidable challenges in training value models for complex problems, the potential benefits of overcoming these difficulties are substantial.
@@
However, training a perfect value model in Long COT tasks presents significant challenges. First, learning a low-bias value model is non-trivial given the long trajectory and the instability of learning value in a bootstrapped way. Second, handling both short and long responses simultaneously is also challenging, as they might exhibit very distinct preferences towards the bias-variance trade-off during optimization. Last but not least, the sparsity of the reward signal from verifiers is further exacerbated by the long CoT pattern, which intrinsically requires better mechanisms to balance exploration and exploitation. To address the aforementioned challenges and fully unleash the potential of value-based methods in reasoning tasks, we present Value Augmented proximal Policy Optimization (VAPO), a value-based RL training framework. VAPO draws inspiration from prior research works such as VC-PPO [30] and DAPO [29], and further extends their concepts.
@@
    We introduce VAPO, the first value-based RL training framework to outperform value-free methods on long COT tasks significantly. VAPO not only demonstrates remarkable superiority in terms of performance but also showcases enhanced training efficiency, streamlining the learning process and underscoring its potential as a new benchmark in the field.
@@
VAPO is an effective reinforcement learning system that brings together these improvements. These enhancements work together smoothly, leading to a combined result that’s better than the sum of the individual parts. We conduct experiments using the Qwen2.5-32B pre-trained model, ensuring no SFT data is introduced in any of the experiments, to maintain comparability with related works (DAPO and DeepSeek-R1-Zero-Qwen-32B). The performance of VAPO improves from vanilla PPO a score of 5 to 60, surpassing the previous SOTA value-free methods DAPO [29] by 10 points. More importantly, VAPO is highly stable — we don’t observe any crashes during training, and the results across multiple runs are consistently similar.
@@ 4.1 Mitigating Value Model Bias over Long Sequences
Building upon the analysis of value-based models presented in section 3.1, we propose to use Value-Pretraining and decoupled-GAE to address the critical challenges in value model bias over long sequences. Both of these two techniques draw upon methodologies previously introduced in VC-PPO.
@@ 6 Related Work
OpenAI O1 [16] introduces a profound paradigm shift in LLMs, characterized by extended reasoning before delivering a final response [28, 19, 5]. DeepSeek R1 [6] open-sources both its training algorithm (the value-free GRPO [22]) and its model weights, which are comparable in performance to O1. DAPO [29] identifies previously undisclosed challenges such as entropy collapse encountered during the scaling of value-free LLM RL, and proposes four effective techniques to overcome these challenges, achieving SOTA industry-level performance. Recently, Dr. GRPO [12] removes both the length and std normalization terms in GRPO. On the other hand, ORZ [9] follows PPO and utilizes a value model for advantage estimation, proposing Monte Carlo estimation instead of Generalized Advantage Estimation. However, they could just achieves a comparable performance to value-free method like GRPO and DAPO. In this paper, we also follow the value-model approach and propose VAPO, which outperforms the SOTA value-free algorithm DAPO.
@@ 7 Conclusion
In this paper, we propose an algorithm named VAPO, which leveraging the Qwen2.5-32B model, achieves the SOTA performance on the AIME24 benchmark. By introducing seven novel techniques atop PPO, which focus on refining value learning and balancing exploration, our value-based approach outperforms contemporary value-free methods like GRPO and DAPO. The work provides a robust framework for advancing large language models in reasoning-intensive tasks.

### v3

@@ Abstract
We present VAPO, Value-model-based Augmented Proximal Policy Optimization framework for reasoning models., a novel framework tailored for reasoning models within the value-model-based paradigm. Benchmarked the AIME 2024 dataset, VAPO, built on the Qwen 32B pre-trained model, attains a state-of-the-art score of $\mathbf{60.4}$. In direct comparison under identical experimental settings, VAPO outperforms the previously reported results of DeepSeek-R1-Zero-Qwen-32B and DAPO by more than 10 points. The training process of VAPO stands out for its stability and efficiency. It reaches state-of-the-art performance within a mere 5,000 steps. Moreover, across multiple independent runs, no training crashes occur, underscoring its reliability. This research delves into long chain-of-thought (long-CoT) reasoning using a value-model-based reinforcement learning framework. We pinpoint three key challenges that plague value-model-based methods: value model bias, the presence of heterogeneous sequence lengths, and the sparsity of reward signals. Through systematic design, VAPO offers an integrated solution that effectively alleviates these challenges, enabling enhanced performance in long-CoT reasoning tasks.
@@ 1 Introduction
In the Large Language Models (LLM) [3, 15, 11, 28, 25, 2, 4] RL training, value-model-free methods like GRPO [22] and DAPO [29] have demonstrated remarkable effectiveness. These approaches eliminate the computational overhead of learning a value model, instead computing advantage solely based on the final reward of the entire trajectory. The trajectory-level advantage is then directly assigned as the token-level advantage for each position in the sequence. When training a reliable value model is particularly challenging, value-model-free methods deliver an accurate and stable baseline for advantage calculation by averaging the rewards across multiple trajectories within a group. This group-based reward aggregation mitigates the need for explicit value estimation, which often suffers from instability in complex tasks. Consequently, value-model-free methods have gained significant traction in addressing difficult problems such as long-CoT reasoning, with substantial research efforts focused on optimizing their frameworks.
@@
Despite the notable success achieved by the value-model-free methods, we argue that value-model-based approaches possess a higher performance ceiling if the challenges in training value models can be addressed. First, value models enable more precise credit assignment by accurately tracing the impact of each action on subsequent returns, facilitating finer-grained optimization [21]. This is particularly critical for complex reasoning tasks, where subtle errors in individual steps often lead to catastrophic failures, and it remains challenging for model optimizing under value-model-free frameworks [30]. Secondly, in contrast to the advantage estimates derived from Monte Carlo methods in value-model-free approaches, value models can provide lower-variance value estimates for each token, thereby enhancing training stability. Furthermore, a well-trained value model exhibits inherent generalization capabilities, enabling more efficient utilization of samples encountered during online exploration. This significantly elevates the optimization ceiling of reinforcement learning algorithms. Consequently, despite the formidable challenges in training value models for complex problems, the potential benefits of overcoming these difficulties are substantial.
@@
However, training a perfect value model in Long COT tasks presents significant challenges. First, learning a low-bias value model is non-trivial given the long trajectory and the instability of learning value in a bootstrapped way. Second, handling both short and long responses simultaneously is also challenging, as they might exhibit very distinct preferences towards the bias-variance trade-off during optimization. Last but not least, the sparsity of the reward signal from verifiers is further exacerbated by the long CoT pattern, which intrinsically requires better mechanisms to balance exploration and exploitation. To address the aforementioned challenges and fully unleash the potential of value-model-based methods in reasoning tasks, we present Value Augmented proximal Policy Optimization (VAPO), a value-model-based RL training framework. VAPO draws inspiration from prior research works such as VC-PPO [30] and DAPO [29], and further extends their concepts.
@@
    We introduce VAPO, the first value-model-based RL training framework to outperform value-model-free methods on long COT tasks significantly. VAPO not only demonstrates remarkable superiority in terms of performance but also showcases enhanced training efficiency, streamlining the learning process and underscoring its potential as a new benchmark in the field.
@@
VAPO is an effective reinforcement learning system that brings together these improvements. These enhancements work together smoothly, leading to a combined result that’s better than the sum of the individual parts. We conduct experiments using the Qwen2.5-32B pre-trained model, ensuring no SFT data is introduced in any of the experiments, to maintain comparability with related works (DAPO and DeepSeek-R1-Zero-Qwen-32B). The performance of VAPO improves from vanilla PPO a score of 5 to 60, surpassing the previous SOTA value-model-free methods DAPO [29] by 10 points. More importantly, VAPO is highly stable — we don’t observe any crashes during training, and the results across multiple runs are consistently similar.
@@ 4.1 Mitigating Value Model Bias over Long Sequences
Building upon the analysis of value-model-based models presented in section 3.1, we propose to use Value-Pretraining and decoupled-GAE to address the critical challenges in value model bias over long sequences. Both of these two techniques draw upon methodologies previously introduced in VC-PPO.
@@ 6 Related Work
OpenAI O1 [16] introduces a profound paradigm shift in LLMs, characterized by extended reasoning before delivering a final response [28, 19, 5]. DeepSeek R1 [6] open-sources both its training algorithm (the value-model-free GRPO [22]) and its model weights, which are comparable in performance to O1. DAPO [29] identifies previously undisclosed challenges such as entropy collapse encountered during the scaling of value-model-free LLM RL, and proposes four effective techniques to overcome these challenges, achieving SOTA industry-level performance. Recently, Dr. GRPO [12] removes both the length and std normalization terms in GRPO. On the other hand, ORZ [9] follows PPO and utilizes a value model for advantage estimation, proposing Monte Carlo estimation instead of Generalized Advantage Estimation. However, they could just achieves a comparable performance to value-model-free method like GRPO and DAPO. In this paper, we also follow the value-model-based approach and propose VAPO, which outperforms the SOTA value-model-free algorithm DAPO.
@@ 7 Conclusion
In this paper, we propose an algorithm named VAPO, which leveraging the Qwen2.5-32B model, achieves the SOTA performance on the AIME24 benchmark. By introducing seven novel techniques atop PPO, which focus on refining value learning and balancing exploration, our value-model-based approach outperforms contemporary value-model-free methods like GRPO and DAPO. The work provides a robust framework for advancing large language models in reasoning-intensive tasks.